Present and Future Value of an Uneven Cash Flow Stream

An investment will pay $200 at the end of each of the next 3 years, $400 at the end of Year 4, $500 at the end of Year 5, and $600 at the end of Year 6. If other investments of equal risk earn 9% annually, what is this investment's present value? Its future value? Do not round intermediate calculations. Round your answers to the nearest cent.

Present value: $  

Future value: $  
Answer #1

Present value=Cash flows*Present value of discounting factor(rate%,time period)

=200/1.09+200/1.09^2+200/1.09^3+400/1.09^4+500/1.09^5+600/1.09^6

=$1472.36(Approx).

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

Future value=1472.36*(1.09)^6

=$2469.29(Approx).